OpenStudy (anonymous):

Evaluate the integral for all positive whole numbers n: \[ \int^1_0 \left(\log\frac{1}{x}\right)^n dx \]

OpenStudy (zarkon):

n!

OpenStudy (dumbcow):

wow how did you get n! ? using substitution i seem to get a repeating integration by parts u = log(x) --> x = e^u du = dx/x \[ - \int\limits_{0}^{1}e^{u}*u^{n} du

OpenStudy (ybarrap):

for positive whole numbers n plus zero http://www.wolframalpha.com/input/?i=integrate+%28+log+1%2Fx+%29^n+dx+from+x%3D0+to+1

OpenStudy (zarkon):

integration by parts

OpenStudy (zarkon):

\[u=\left(\ln(1/x)\right)^n\] \[du=n(\ln(1/x))^{n-1}\frac{1}{1/x}\frac{-1}{x^2}=-n(\ln(1/x))^{n-1}(1/x)\] \[dv=dx\] \[v=x\]

OpenStudy (zarkon):

forgot a dx

OpenStudy (zarkon):

\[du=n(\ln(1/x))^{n-1}\frac{1}{1/x}\frac{-1}{x^2}dx=-n(\ln(1/x))^{n-1}(1/x)dx\]

OpenStudy (anonymous):

You can also use substitution to show that it equals \[ \Gamma(n+1) \]

OpenStudy (zarkon):

then use the formula for parts....you will get the reduction formula

OpenStudy (zarkon):

both just as trivial